### v2

@@ 3.1.3 Illumination adjustment mechanism
Photographic images in real-world scenarios often exhibit shadows due to lighting angles, with shadow locations varying based on illumination direction. Most shadows originate from camera equipment, with the darkest areas typically localized near one of the four image corners. We proposed the illumination adjustment mechanism, which first calculates the overall image brightness using Equation 3, derived from the standard-definition television standard ITU-R BT.601[28]:
@@
The data generator integrates these three operations (wave distortion, cylindrical warping, and illumination adjustment) through coordinated parameter randomization. Parameter constraints imitate natural deformations: for instance, $A$ and $\omega$ in Equation 1 maintain an inversely proportional relationship, while $F$ and $c$ in Equation 2 are constrained to avoid extreme values. These inter-parameter dependencies guide the system toward generating physically plausible image distortions.
@@ 3.2 Derived dataset DWTAL
The data generator enables large-scale deformed tabular data synthesis from limited mildly deformed tables. While the dataset incorporates 150 self-collected samples, it primarily leverages two public natural scene datasets (TAL-OCR and WTW). Initially, mildly deformed tables from TAL-OCR and WTW are selected to generate fine-grained segmentation masks using corner coordinates. These masks are then iteratively processed by the data generator to produce diverse deformed table images. The two source datasets exhibit distinct characteristics: TAL-OCR contains education-oriented images with clear table structures and simplified backgrounds, leading to relatively simpler derived data. To enhance diversity, self-collected samples are merged with TAL-OCR-derived data to form a compact dataset named DWTAL-s. In contrast, WTW features more complex backgrounds and richer content, enabling the creation of a larger, more challenging dataset called DWTAL-l.
Dataset Splitting. Both datasets follow identical partitioning protocols. To ensure uniform distribution of deformation types across training and test sets, 80% of each dataset is randomly allocated for training and 20% for testing. Final counts reach 7,012 training and 1,753 test images in DWTAL-s, versus 15,616 training and 3,904 test images in DWTAL-l.
The derived datasets inherit core features from their parent collections, including curved deformations, perspective distortions, multicolor backgrounds, surface irregularities, and illumination variations. A key enhancement lies in the pixel-level instance segmentation annotations for table cells, though all images exclusively contain single-table instances. Additionally, a dataset version with logical coordinate annotations will be publicly released to support broader research applications.
WTW
TAL-OCR
Collection
@@ 4.1 Overall architecture
Figure 3 illustrates the overall architecture of OG-HFYOLO, which adopts a YOLO-like framework divided into three core components: a feature extraction backbone, a feature fusion neck, and a detection head. To enhance texture feature extraction, the model first feeds the initial downsampled feature map into a Gradient Orientation-aware Extractor (GOE), generating feature maps with enriched texture information. The backbone employs CSP-Unit modules, each comprising a $3\times 3$ downsampling convolution, SiLU activation, batch normalization, and a Cross-Stage Partial Network (CSP) [29] used by YOLOv5. Through five downsampling operations, the backbone produces multi-scale feature maps P3, P4, P5.
During feature fusion, the extracted features undergo integration via the FPN-PAN [30] pathway. Distinct from standard YOLO implementations, our model incorporates Heterogeneous Kernel Cross Fusion (HKCF) module after skip connections to enhance cross-scale objective feature interactions. The fused features are then processed through CSP blocks for final refinement before being delivered to the detection head. The detection head retains YOLOv5’s anchor-based [31] design, performing classification and bounding box regression using predefined anchor boxes, with Non-Maximum Suppression [32] (NMS) filtering redundant detections.
@@ 4.2 Gradient Orientation-aware Extractor
We consider object density the primary challenge in spatial coordinate localization for deformed table cells. Inaccurate contour extraction by the model risks merging adjacent small cells into erroneously detected large ones. This aggregation progressively obscures scale variations during feature fusion, ultimately degrading prediction accuracy.
The accuracy of instance segmentation in target-intensive scenarios is limited by the model’s ability to perceive contour details, and the essential challenge lies in how to effectively model the gradient strength and orientation sensitivity of texture regions. The Histogram of Oriented Gradients (HOG) [33] offers critical insights here: As shown in Figure 4, HOG divides the image into multiple grids, decouples gradient features into gradient direction and gradient intensity within each grid, and aggregates these attributes into orientation-specific histograms. Its core philosophy lies in simultaneously capturing gradient magnitude and directional priors to enhance the geometric discriminability of contours, a principle that directly informs our approach to addressing contour ambiguity in dense cell detection.
Figure 4: HOG feature extraction core schematic diagram: the left inset demonstrates gradient orientation features within the red grid, while the right inset illustrates gradient magnitude features within the same grid. These two key features allow for comprehensive features of edge texture information.
The GOE adopts the same philosophy, focusing on enabling the model to learn both gradient intensity and gradient direction features of contour details, which effectively enhances its recognition capability for targets with complex contours and dense arrangements.
@@
Figure 5 illustrates the internal architecture of the proposed Gradient Orientation-aware Extractor (GOE). The module takes a feature map encoding texture information as input. The GOE first employs distinct decoupling operators (Equation 5) to decompose the input feature map $\mathbf{I}$ into a horizontal gradient orientation feature map ${G}_{x}$ and a vertical gradient orientation feature map ${G}_{y}$. To preserve the directional priors of the decoupling operators $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$, their weights are initialized as conventional edge operators. During training, kernel parameter constraints are relaxed, allowing the network to adaptively adjust gradient response weights based on task requirements (e.g., deformation edge enhancement, noise suppression), thereby overcoming the geometric limitations inherent to standard edge operators.
@@
Orientation Attention $\Phi$: The orientation attention $\Phi$ in the GOE module is fundamentally a learnable convolutional operation. The module initializes the convolutional kernel weights $W$ by emulating the orientation binning technique used in Histogram of Oriented Gradients (HOG). Specifically, the continuous angular space $[0,\pi)$ is uniformly discretized into n orientation bins, with the central angle of each bin defined as $\theta_{i}=\frac{i\pi}{n},(i=0,1,...,n-1)$, For each orientation bin $\theta_{i}$, the corresponding unit directional basis vectors in Cartesian coordinates are computed via the polar coordinate transformation formula:$[e_{x}^{i}=cos(\theta_{i}),e_{y}^{i}=sin(\theta_{i})]$. These basis vectors are then concatenated and expanded across two dimensions to construct the initialized kernel weights $W\in R^{n\times 2\times 1\times 1}$.
The hierarchical feature learning mechanism of convolutional neural networks demonstrates that shallow high-resolution feature maps exhibit stronger capabilities in capturing low-level visual features such as edges and textures, while deeper layers focus more on semantic abstraction and global structure modeling[36]. Based on this, we embed the Gradient Orientation Extractor (GOE) after the first convolutional downsampling layer in the backbone network, leveraging its high spatial resolution characteristics to precisely capture directional gradient features. Figure 6 illustrates the effectiveness of the GOE module in early-stage feature extraction from the second downsampled feature map. By incorporating the proposed extractor, the network injects geometric priors at early stages for images with chromatic interference and blurred cell contours, which establishes a fine-grained foundation for subsequent cross-layer feature fusion.
Image
no-GOE
GOE
Figure 6: GOE module effect demonstration.
@@ 4.3 Heterogeneous Kernel Cross Fusion
In addition to dense objective distribution, table cell segmentation also faces challenges arising from extreme scale diversity caused by merged cells: horizontal merging creates wide-spanning targets across multiple columns, while vertical merging generates tall and narrow targets spanning multiple rows. Such objectives require the model to possess multi-granularity perception capabilities. The YOLO series addresses scale variation through multi-scale detection heads, where high-resolution feature maps detect large objectives and low-resolution ones focus on small objectives. However, traditional fixed-size convolutional kernels struggle to adapt to the morphological diversity of cell features in tables. Inspired by YOLO-MS[10], which proposes matching objective diversity through kernel diversity and introduces the Heterogeneous Kernel Selection Protocol[10] (HKSP), we adopt the HKSP concept and integrate asymmetric cross-convolution[37] to design a Heterogeneous Kernel Cross-Fusion (HKCF) module. As shown in Figure 7, the module employs a bottleneck structure to reduce computational complexity. The input feature map $I_{in}$ is first channel-reduced via $1\times 1$ convolution to obtain low-dimensional features $F_{in}$. The HKCF module $HKCF^{(k)}$ with kernel size $k$ then extracts features $F_{out}$ in this reduced channel space. Finally, $F_{out}$ and $F_{in}$ are concatenated along the channel dimension, and the original channel dimension is restored through another $1\times 1$ convolution to output the feature map $I_{out}$. The computational flow can be formulated as:
@@
As shown in Equation 10, within the Heterogeneous Kernel Cross-Fusion module $HKCF^{(k)}$, the input feature map $F_{in}$ first undergoes processing through the Channel Attention Bridge[38] (CAB). The CAB employs global average pooling and a multilayer perceptron to generate channel-wise weights, dynamically enhancing feature channels strongly correlated with target morphology, thereby alleviating information loss caused by the bottleneck structure’s dimensionality reduction. To address the extreme narrow-tall and wide-flat challenges posed by merged cells, the module replaces standard convolution with heterogeneous cross convolution $HKCF^{(k)}$. This operation deploys horizontally extended kernels ($1\times k$) and vertically extended kernels ($k\times 1$) in parallel. These dual branches respectively capture horizontal continuity features for column-spanning merged cells and vertical long-range dependencies for row-spanning merged cells. The outputs of these two branches sum to precisely adapt to the geometric characteristics of targets at varying scales. Following HKSP, cross-convolutions with kernel sizes 3, 5, 7 are progressively applied from shallow to deep feature fusion maps. Finally, the original input $F_{in}$ is added to the output feature map from $HXConv^{(k)}$ by residual connection, producing the final output feature map $F_{out}$.
@@ 4.4 Loss function optimization
In anchor-based instance segmentation frameworks, the object loss $\mathcal{L}_{obj}$ typically relies on Intersection over Union (IoU) and its variants to measure geometric deviations between predicted and ground-truth bounding boxes. However, the conventional CIoU loss[39] suffers from ambiguous optimization directions for extreme aspect ratio targets due to its coupled computation of aspect ratio and center distance, leading to reduced bounding box regression accuracy. To address this, we replace CIoU with the EIoU loss[40], which explicitly decouples width-height optimization paths, enabling targeted gradient direction adjustment for aspect ratio-sensitive targets.
[…]
@@
where $N$ denotes the total number of pixels in the mask, $\mathbf{M}_{\text{gt}}$ represents the ground truth mask, and $\mathbf{M}_{\text{pred}}$ is the predicted mask.
@@
where $A$ indicates the area (pixel count) of the target instance for normalization, while $H^{\prime}$ and $W^{\prime}$ represent the height and width of the cropped region $crop(\Omega)$, respectively, which can be spatially averaged for loss computation, analogous to global average pooling.
However, relying solely on the binary cross-entropy loss $\mathcal{L}_{BCE}$ neglects the inherent similarity between cell backgrounds and the overall table background, where background pixels dominate in quantity. To prevent the model from biasing towards the high-frequency background class, we integrate Dice Loss[41] into the base segmentation loss $\mathcal{L}_{base}$. Unlike cross-entropy loss, which focuses on per-pixel probability calibration, Dice Loss optimizes for region overlap, prioritizing the structural integrity of segmented targets. This approach enhances edge alignment and regional continuity. The Dice Loss formula is defined in Equation 13, where $\odot$ denotes element-wise multiplication.
$\mathcal{L}_{\text{Dice}}=1-\frac{2\sum\mathbf{M}_{\text{pred}}\odot\mathbf{M}_{\text{gt}}}{\sum\mathbf{M}_{\text{pred}}+\sum\mathbf{M}_{\text{gt}}}$ (13)
The base loss $\mathcal{L}_{base}$ is formed by summing the two losses according to Equation 14. In the original loss function, the inverse proportional function for normalization aims to eliminate the influence of target size on loss weighting. While this effectively prevents training bias toward large-target optimization, it adversely affects the optimization of small targets. Specifically, the inverse proportional function applies overly aggressive weighting to losses from small-area targets. To address this issue, this paper reconsiders the allocation of dynamic weights: to avoid excessive weighting amplification for small targets, the proposed dynamic weighting scheme incorporates a smoother logarithmic function on top of the original normalization weights. The calculation is as follows: 15
$\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{BCE}}+\mathcal{L}_{\text{Dice}}$ (14)
$W=1+\log\left(\frac{1}{A}\right)$ (15)
Integrating the above considerations, we obtain the scale-aware loss function $\mathcal{L}_{{scale-aware}}$, as defined in Equation 16, where $n$ denotes the number of instances in the image, $H_{i}^{\prime}$ and $W_{i}^{\prime}$ represent the height and width of the cropped region for the instance $i$, and $A_{i}$ corresponds to the area of the instance $i$.
@@
$\mathcal{L}_{{scale-aware}}$ $=\frac{1}{n}\sum_{i=1}^{n}\left[W_{i}\frac{1}{A_{i}}\sum_{(x,y)\in{crop}(\Omega)}\left[\frac{1}{H_{i}^{\prime}W_{i}^{\prime}}\mathcal{L}_{{base}}(x,y)\right]\right]$ (16)
$=\frac{1}{n}\sum_{i=1}^{n}\left[\left(1+\log\left(\frac{1}{A_{i}}\right)\right)\cdot\frac{1}{A_{i}}\sum_{(x,y)\in{crop}(\Omega)}\left[\frac{1}{H_{i}^{\prime}W_{i}^{\prime}}\left(\mathcal{L}_{{BCE}}+\mathcal{L}_{{Dice}}\right)\right]\right]$

### v3

@@ 3.1.3 Illumination adjustment mechanism
Photographic images in real-world scenarios often exhibit shadows due to lighting angles, with shadow locations varying with illumination direction. Most shadows originate from camera equipment, with the darkest areas typically localized near one of the four image corners. We propose an illumination adjustment mechanism that first calculates overall image brightness using Equation 3. This equation is derived from the standard-definition television standard ITU-R BT.601 [28]:
@@
The data generator integrates three operations (wave distortion, cylindrical warping, and illumination adjustment) through coordinated parameter randomization. For wave distortion, the amplitude $A$ is uniformly sampled from $[10,50]$, while the wavelength $\omega$ is dynamically constrained to avoid unrealistic high-frequency distortions: ${\omega}{\text{min}}$ increases proportionally to $A$ via a scaling factor $s\in[1,5]$, and $\omega$ is finally selected from $[{\omega}{\text{min}},800]$. For cylindrical warping, the curvature axis $c$ follows a truncated normal distribution $\mathcal{N}(2,0.7^{2})$ within $[1,5]$, and the deformation factor $F$ is adaptively sampled from a range determined by $c$—smaller $c$ (near the image center) corresponds to milder distortions ($F\in[0.7,0.85]$), while larger $c$ (offset from the center) reduces $F_{\text{min}}$ linearly to enforce physically plausible attenuation. For illumination adjustment, shadows are added only if the image brightness exceeds a threshold, with center brightness $cb\in[0.6,0.9]$ and edge brightness $eb\in[0.1,0.3]$ controlling the intensity gradient. These inter-parameter dependencies ensure that generated distortions mimic natural variations while avoiding implausible artifacts.
@@ 3.2 Derived dataset DWTAL
The data generator enables large-scale synthesis of deformed table data from limited mildly deformed tables. Although DWTAL datasets includes 150 collected samples, it primarily relies on two public natural scene datasets: TAL-OCR and WTW. First, mildly deformed tables from TAL-OCR and WTW are selected to generate fine-grained segmentation masks using corner coordinates. These masks are then iteratively processed by the data generator to produce diverse deformed table images. The two source datasets have distinct characteristics. TAL-OCR contains education-focused images with clear table structures and simplified backgrounds, resulting in relatively simpler derived data. To enhance diversity, collected samples are merged with TAL-OCR-derived data to form a compact dataset called DWTAL-s, which has a total of 8,765 images. In contrast, WTW has more complex backgrounds and richer content, enabling the creation of a larger, more challenging dataset named DWTAL-l, which has a total of 19,520 images.
Dataset Splitting. Both datasets follow identical partitioning protocols. To ensure uniform distribution of deformation types across training and test sets, 80% of each dataset is randomly allocated for training and 20% for testing. Finally, the DWTAL-s dataset has 7,012 images in the training set and 1,753 images in the test set, while the DWTAL-l dataset has 15,616 images in the training set and 3,904 images in the test set.
The derived datasets retain key characteristics from their parent collections, such as curved deformations, perspective distortions, multicolor backgrounds, surface irregularities, and illumination variations. A significant improvement is the inclusion of pixel-level instance segmentation annotations for table cells. However, all images contain only single-table instances. Furthermore, a dataset version with logical coordinate annotations has been publicly released to facilitate wider research applications.
@@ 4.1 Overall architecture
Figure 3 illustrates the overall architecture of OG-HFYOLO, which adopts a YOLO framework and comprises three core components: the feature extraction backbone, the feature fusion neck, and the detection head. To enhance texture feature extraction, the model feeds the downsampled feature maps into the Gradient Orientation-aware Extractor (GOE) and concatenates them with the original input features along the channel dimension, generating fused feature representations enriched with texture information. The backbone employs CSP-Unit modules. Each module includes a $3\times 3$ downsampling convolution, SiLU activation, batch normalization, and the Cross-Stage Partial Network (CSP) [29] utilized in YOLOv5. Through five downsampling operations, the backbone produces multi-scale feature maps P3, P4, and P5.
During feature fusion, the extracted features are integrated via the FPN-PAN [30] pathway. Unlike standard YOLO implementations, our model incorporates a Heterogeneous Kernel Cross Fusion (HKCF) module after skip connections to enhance cross-scale feature interactions. The fused features are then refined via CSP blocks before being passed to the detection head. The fused features are then refined via CSP blocks before being passed to the detection head. The detection head retains YOLOv5’s anchor-based [31] design, performing classification and bounding box regression using predefined anchor boxes, with Non-Maximum Suppression [32] (NMS) filtering redundant detections.
@@ 4.2 Gradient Orientation-aware Extractor
We identify object density as the primary challenge in spatial coordinate localization for deformed table cells. Inaccurate contour extraction by the model may lead to the merging of adjacent small cells into erroneously detected large ones. This aggregation progressively obscures scale variations during feature fusion, ultimately reducing prediction accuracy.
The instance segmentation accuracy in densely packed scenes is limited by the model’s ability to perceive contour details. The Histogram of Oriented Gradients (HOG) [33] is designed to detect complex human contours, positing that fine gradients and precise orientation binning are crucial for capturing intricate outlines. Its core principle provides significant insights for dense and complex object detection: as shown in Figure 4, HOG divides the image into multiple grids and decouples the gradient features within each grid into gradient direction and gradient intensity. It then statistically aggregates the gradient intensities in each direction to form a histogram. The core idea is to simultaneously capture the magnitude information of the gradients and the directional priors, thereby enhancing the geometric discriminability of the contours.
Figure 4: HOG feature extraction core schematic diagram: the left inset demonstrates gradient orientation features within the red grid, while the right(Heatmap: The brighter the area, the higher the intensity) inset illustrates gradient magnitude features within the same grid. These two key features allow for comprehensive features of edge texture information.
The Gradient Orientation-aware Extractor(GOE) operates on this principle, enabling the model to learn both gradient intensity and gradient direction features of contour details. This effectively enhances its recognition capability for objects with complex contours and dense arrangements.
@@
Figure 5 illustrates the internal architecture of the proposed Gradient Orientation-aware Extractor (GOE). The module takes a feature map encoding texture information as input. The GOE first uses distinct decoupling operators (Equation 5) to decompose the input feature map $\mathbf{I}$ into a horizontal gradient orientation feature map ${G}_{x}$ and a vertical gradient orientation feature map ${G}_{y}$.
@@
To preserve the directional priors of the decoupling operators $\mathcal{G}_{x}$ and $\mathcal{G}_{y}$, their weights are initialized as conventional edge operators. During training, convolution kernel parameter constraints are relaxed. This allows the network to adaptively adjust gradient response weights according to task requirements (e.g., deformation edge enhancement, noise suppression), thus overcoming the inherent geometric limitations of standard edge operators.
@@
Orientation Attention $\Phi$: The orientation attention mechanism $\Phi$ in the Gradient Orientation-aware Extractor (GOE) module is designed to encode gradient directionality through a learnable convolutional operation, combining geometric priors with adaptive feature learning. Inspired by the orientation binning strategy in Histogram of Oriented Gradients (HOG), we initialize the convolutional kernels using directional basis vectors derived from polar coordinate transformations. Specifically, the continuous angular space $[0,\pi)$ is uniformly divided into $n$ discrete bins, with the central angle of the $i$-th bin defined as $\theta_{i}=\frac{i\pi}{n}$. For each angle $\theta_{i}$, the corresponding Cartesian unit vector is computed as:
$\mathbf{e}_{i}=\begin{bmatrix}\cos\theta_{i}\\ \sin\theta_{i}\end{bmatrix}\in\mathbb{R}^{2},i=0,1,2,\ldots,n-1.$ (9)
explicitly representing gradient orientation in orthogonal coordinates. These vectors are structured into a parameter matrix $W\in\mathbb{R}^{n\times 2\times 1\times 1}$, where each row corresponds to a directional basis:
$W=\begin{bmatrix}\cos\theta_{0}&\sin\theta_{0}\\ \cos\theta_{1}&\sin\theta_{1}\\ \vdots&\vdots\\ \cos\theta_{n-1}&\sin\theta_{n-1}\end{bmatrix}\otimes\mathbf{I}_{1\times 1},$ (10)
with $\otimes$ denoting the outer product expansion to match the 4D tensor structure of convolutional kernels. This matrix serves as the initial weights of a $1\times 1$ convolution layer, mapping the input gradient fields (horizontal and vertical components) to $n$ orientation-specific feature channels. During training, the kernel weights are dynamically optimized through backpropagation, allowing the network to suppress noisy orientations while amplifying discriminative directional patterns. Unlike static HOG descriptors, this learnable mechanism adapts to task-specific requirements—for example, enhancing deformation boundaries in densely packed table cells. The final output combines orientation-aware features with the original texture representations through channel-wise concatenation, providing a geometrically grounded yet flexible foundation for downstream segmentation tasks.
The hierarchical feature learning mechanism in convolutional neural networks shows that shallow, high-resolution feature maps are better at capturing low-level visual features such as edges and textures. In contrast, deeper layers focus more on semantic abstraction and global structure modeling [36]. Based on this, we integrate the Gradient Orientation Extractor (GOE) after the first convolutional downsampling layer in the backbone network. This leverages the high spatial resolution of shallow layers to precisely capture directional gradient features. Figure 6 demonstrates the effectiveness of the GOE module in early feature extraction from the second downsampled feature map. By incorporating the proposed extractor, the network introduces geometric priors in the early stages for images with chromatic interference and blurred cell contours. This establishes a fine-grained foundation for subsequent cross-layer feature fusion.
Figure 6: GOE module effect demonstration: The second line shows the effect of the second downsampling feature map of the model in the backbone without adding the GOE module, and the third line shows the effect of the second downsampling feature map of the model in the backbone with adding the GOE module.
@@ 4.3 Heterogeneous Kernel Cross Fusion
In addition to dense objective distribution, table cell segmentation also faces challenges due to extreme scale diversity caused by merged cells. Horizontal merging creates wide-spanning objects across multiple columns, while vertical merging generates tall and narrow objects spanning multiple rows. Such objectives demand multi-granularity perception capabilities from the model. The YOLO series addresses scale variation through multi-scale detection heads. High-resolution feature maps detect small objectives, while low-resolution ones focus on large objectives. However, traditional fixed-size convolutional kernels struggle to adapt to the morphological diversity of cell features in tables. Inspired by YOLO-MS [10], which proposes matching objective diversity through kernel diversity and introduces the Heterogeneous Kernel Selection Protocol (HKSP) [10], we incorporate the HKSP concept and integrate asymmetric cross-convolution [37] to design a Heterogeneous Kernel Cross-Fusion (HKCF) module. As shown in Figure 7, the module uses a bottleneck structure to reduce computational complexity. As shown in Figure 7, the module uses a bottleneck structure to reduce computational complexity. The input feature map $I_{in}$ is first channel-reduced using $1\times 1$ convolution to obtain low-dimensional features $F_{in}$. The $HKCF^{(k)}$ module with kernel size $k$ then extracts features $F_{out}$ in this reduced channel space. Finally, $F_{out}$ and $F_{in}$ are concatenated along the channel dimension. The original channel dimension is restored via another $1\times 1$ convolution to output the feature map $I_{out}$. The computational flow can be formulated as:
@@
As shown in Equation 12, within the Heterogeneous Kernel Cross-Fusion module $HKCF^{(k)}$, the input feature map $F_{in}$ first undergoes processing through the Channel Attention Bridge[38] (CAB). The CAB uses global average pooling and a multilayer perceptron to generate channel-wise weights, dynamically enhancing feature channels strongly correlated with object morphology, thereby alleviating information loss caused by the bottleneck structure’s dimensionality reduction. To address the extreme narrow-tall and wide-flat challenges posed by merged cells, the module replaces standard convolution with heterogeneous cross convolution $HKCF^{(k)}$. This operation deploys horizontally extended kernels ($1\times k$) and vertically extended kernels ($k\times 1$) in parallel. These dual branches respectively capture horizontal continuity features for column-spanning merged cells and vertical long-range dependencies for row-spanning merged cells. The outputs of these two branches sum to precisely adapt to the geometric characteristics of objects at varying scales. Following HKSP, cross-convolutions with kernel sizes 3, 5, 7 are progressively applied from shallow to deep feature fusion maps. Finally, the original input $F_{in}$ is added to the output feature map from $HXConv^{(k)}$ by residual connection, producing the final output feature map $F_{out}$.
@@ 4.4 Loss function optimization
In anchor-based instance segmentation frameworks, the object loss $\mathcal{L}_{obj}$ typically depends on Intersection over Union (IoU) and its variants to measure geometric deviations between predicted and ground-truth bounding boxes. However, the conventional CIoU loss [39] has ambiguous optimization directions for objects with extreme aspect ratios because it couples aspect ratio and center distance computations. This results in reduced bounding box regression accuracy. To address this issue, we replace CIoU with the EIoU loss [40]. EIoU explicitly decouples width-height optimization paths, enabling objected gradient direction adjustments for objects sensitive to aspect ratios.
For mask loss $\mathcal{L}_{mask}$,the YOLO framework uses the binary cross-entropy loss ($\mathcal{L}_{BCE}$) as follows:
@@
where $N$ denotes the total number of pixels in the mask, $\mathbf{M}^{i}_{\text{gt}}$ represents the i-th pixel in the true mask, and $\mathbf{M}^{i}_{\text{pred}}$ is the i-th pixel in the predicted mask.
However, relying solely on the binary cross-entropy loss $\mathcal{L}_{BCE}$ neglects the inherent similarity between cell backgrounds and the overall table background, where background pixels dominate in quantity. To prevent the model from biasing towards the high-frequency background class, we integrate Dice Loss[41] into the base segmentation loss $\mathcal{L}_{base}$. Unlike cross-entropy loss, which focuses on per-pixel probability calibration, Dice Loss optimizes for region overlap, prioritizing the structural integrity of segmented objects. This approach enhances edge alignment and regional continuity. The Dice Loss formula is defined in Equation as follows:
$\mathcal{L}_{\text{Dice}(x,y)}=1-\frac{2\sum\mathbf{M}_{\text{pred}}(x,y)\odot\mathbf{M}_{\text{gt}}(x,y)}{\sum\mathbf{M}_{\text{pred}}(x,y)+\sum\mathbf{M}_{\text{gt}}(x,y)}$ (14)
where $\odot$ denotes element-wise multiplication.
The base loss $\mathcal{L}_{base}$ combines binary cross-entropy and Dice loss through summation:
$\mathcal{L}_{\text{base}}=\mathcal{L}_{\text{BCE}}+\mathcal{L}_{\text{Dice}}.$ (15)
In the original loss function, YOLO normalizes the loss to balance the influence of objects of different sizes through inverse of the object’s area term as follows:
@@
where $H^{\prime}$ and $W^{\prime}$ represent the height and width of the bounding-box cropped region $crop(\Omega)$ for the instance, and $A$ corresponds to the area (pixel count) of the object instance mask for normalization.
This design aims to balance loss magnitudes across objects of different scales, preventing larger objects from dominating the optimization direction due to their pixel advantage. However, this inverse-proportional compensation mechanism exhibits inherent mathematical limitations: while the $1/A$ function approaches infinity as $A\rightarrow 0^{+}$, aligning with the intuitive need for stronger compensation of small objects, its second derivative $\frac{d^{2}}{dA^{2}}(1/A)=2/A^{3}$ reveals rapidly increasing curvature characteristics. This mathematical property causes compensation intensity to exhibit near-vertical growth trends when target area falls below a critical threshold (e.g., $A<0.1$), potentially inducing gradient mutation issues.
To establish a smoother scale adaptation mechanism, we introduce a logarithmic weighting function $W_{s}=1+\log(1/A)$ building upon the original normalization. The mathematical superiority of this design manifests in its differential properties: its first derivative $\frac{dW_{s}}{dA}=-1/A$ responds more gradually to small-scale variations compared to $\frac{d}{dA}(1/A)=-1/A^{2}$ of the original function. This moderation proves particularly significant in the critical small-target regime ($A<0.1$). Simultaneously, its second derivative $\frac{d^{2}W_{s}}{dA^{2}}=1/A^{2}$ maintains convexity while demonstrating substantially lower curvature than the $\frac{1}{A}$ function ($2/A^{3}$), thereby avoiding excessively aggressive compensation.
From a function space perspective, these components form a complementary optimization: $\frac{1}{A}$ provides foundational compensation intensity ensuring necessary loss amplification for small objects, while $\log(1/A)$ acts as a modulator suppressing curvature mutation risks through its progressive growth characteristics. This dual composite structure maintains coherent limiting behavior—both functions approach infinity as $A\rightarrow 0^{+}$, while $W_{s}$ converges to 1 as $A\rightarrow 1^{-}$ ensuring larger objects experience no additional suppression. The resulting scale-aware loss function:
@@
$\mathcal{L}_{{scale-aware}}=\frac{1}{n}\sum_{i=1}^{n}\left[\left(1+\log\frac{1}{A_{i}}\right)\frac{1}{A_{i}}\sum_{(x,y)\in{crop}(\Omega)}\left[\frac{1}{H_{i}^{\prime}W_{i}^{\prime}}\mathcal{L}_{{base}}(x,y)\right]\right]$ (17)
where $n$ denotes the number of instances in the image, $H_{i}^{\prime}$ and $W_{i}^{\prime}$ represent the height and width of the bounding-box cropped region $crop(\Omega)$ for instance $i$, and $A_{i}$ corresponds to the area of instance $i$.
The scale-aware loss achieves three-tiered optimization: at the microscopic level, the logarithmic term smooths gradient variations in the small-target regime; at the mesoscopic level, dual-function synergy maintains compensation continuity across transitional regions ($0.1\leq A_{i}\leq 0.5$); at the macroscopic level, optimization stability for larger targets ($A_{i}>0.5$) is preserved.